OpenStudy (anonymous):

The vertex of a parabola is (5, -3) and another point on the parabola is (1,5). Which point is also on the parabola? A. (0, 3) B. (-1, 9) C. (-1, 15) D. (7, 7)

OpenStudy (accessdenied):

Do you know of the vertex form of a parabola?

OpenStudy (anonymous):

No

OpenStudy (accessdenied):

A parabola may be written in the form: y = a(x - h)^2 + k Where (h, k) is our vertex and a is some constant. We can find a by substituting (5, -3) for (h,k) and (1, 5) in for (x,y) to solve a 1-variable equation. Then we could go through each point listed and check whether it results in a true equation when plugged in for (x,y), keeping (5,-3) in for (h,k).

OpenStudy (anonymous):

k imma write it out now

OpenStudy (anonymous):

is it A (0,3)?

OpenStudy (anonymous):

?

OpenStudy (accessdenied):

o.o woah sorry, I didn't receive any updates on this question! Just a moment!

OpenStudy (accessdenied):

I did not get A... So you have: y = a (x - h)^2 + k (h,k) = (5,-3) y = a (x - 5)^2 - 3 (x,y) = (1, 5) 5 = a(1 - 5)^2 - 3 8 = a (-4)^2 = 16a a = 1/2 You got to this point?

OpenStudy (accessdenied):

We would sub this back in with the vertex: y = 1/2 (x - 5)^2 - 3 We then plug (x,y) = (0,3). Notice the result: 3 = 1/2 (0 - 5)^2 - 3 3 = 25/2 - 3 3 = 12.5 - 3 = 9.5 That is not true! So A is NOT correct.

OpenStudy (anonymous):

how about c if thats not it im not sure

OpenStudy (accessdenied):

I got C as well. You plug in x=-1 15 = 1/2 (-1 - 5)^2 - 3 15 = 1/2 * 6^ 2 - 3 15 = 36/2 - 3 15 = 18 - 3 15 = 15 is a true statement.

OpenStudy (anonymous):

thank you @ Access Denied really appreciate it

OpenStudy (accessdenied):

Glad to be able to help. :)